To prepare 100 mL of 25 mM Tris buffer at pH 8.0, calculate the amount of Tris base and Tris·HCl needed to make the buffer. (Hint: find out the p K a of Tris·HCl and use the Henderson - Hassel Balch Equation) Describe how do you want to make the solution, including how much water to add, how to adjust pH, etc.
Answer #1

Pkb of Tris base = 5.924

pH = 14 - (pkb + log(Tris-HCl/Tris base))

no of mol of tris buffer = V*M/1000

                       = 100*25*10^-3/1000

           = 0.0025 mol

   tris buffer = Tris base + Tris-HCl(x)

    0.0025 = 0.0025-x + x

8 = 14- (5.924+log(x/(0.0025-x))

   x = 0.00136

no of mol of tris base must be taken = 0.0025-0.00136 = 0.00114 mol

amount tris base must be taken = 0.00114*121.14 = 0.138 g

no of mol of tris-HCl must be taken = 0.00136 mol

amount tris-HCl must be taken = 0.00136*157.6 = 0.214 g


take 0.138 g of tris base and 0.214 g tris-HCl dissolve in limited water and then add water up to 100 ml solution.